3/7 of 2/5 is equal to
(a) 5/12 (b) 5/35 (c) 1/35 (d) 6/35

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (d) is correct here
3/7 of 2/5 is equal to = (3/7) × (2/5)

= 6/35


Thanks